In a scanning process, the number of misrecorded pieces of information has a Poisson distribution with parameter λ = 4.24. What is the probability that there are 4 or 8 misrecorded pieces of information?

Round your answer to the nearest ten thousandth.
Answer #1

Solution :

Given that λ = 4.24

=> P(x = 4) + P(x = 8) = (e^-4.24 * 4.24^4)/4! + (e^-4.24 * 4.24^8)/8!

= 0.1940 + 0.0373

= 0.2313

=> The probability that there are 4 or 8 misrecorded pieces of information is 0.2313

Formula :-

=> For poisson distribution , P(x , λ) = (e^(-λ) * λ^x)/x!